# Normative feeling: socially patterned affective mechanisms

### v3

@@ 3.8 Model of mood: an affective layer to modulate behaviour
We use mood mechanism to mean the tendency of an individual toward something i.e. psychological inclination/disposition, and it can mean the environment’s power of disposing the individual to certain modes of interaction. Inspired by this we define how mood is affected by stimuli (environment) and how mood in turn affects behaviour as mood mechanism. We take inspiration for our model of mood from the empirically verified model of mood [14, 15]. More specifically, Eldar et al describe mood state as a scalar summary of "how well things are going": an average of recent reward error (positive or negative) over recent experiences in their case. We define it as average energy gain or loss in our case (its our model’s equivalent of reward, see $S_{energy-gain}$ stimuli). We choose this model as it is a reification, capturing the essence of mood cognition instead of focusing on any particular emotion in any particular organism. Those core features being: stimuli (both internal and external) influencing a fluctuating mood state that in turn modulates behavioural and perceptual tendencies. We also add other stimuli as those chosen in [15]. We chose these stimuli based on the sensory information available to our agents that we believe to be important regarding their survival in the simulated environment (both physical and social). For example, hunger, being punished, seeing others violating a norm, seeing others die. This list is not exhaustive and in theory you could add more (hence the dotted lines between $Stimulus_{2}$ and $Stimulus_{n}$). We randomly initialise the importance of these stimuli and the direction of their effect on mood as a way of seeing when evolution (if at all) utilises them. In other words, we let our evolutionary process specify the functional relationship between its senses and its mood and in turn its mood and its behaviour. This functional specification is what we call mood mechanism. This allows us to see different kinds of mood mechanism emerging without presupposing a mechanism. This allows us to explore how different types of social processes e.g. normative process like punishment pattern mood to produce different mood mechanisms.
@@
Now that we have described the overview of how the mood model operates we will describe it in detail. What follows is a formal description of how mood affects behaviour in our model, with the weights (mood mechanism) affecting the direction and strength of this effect. In more detail, we describe how mood is updated based on stimuli and their evolvable weights, how mood updates with time and how mood consequently affects behaviour also through evolvable weights (see Figure 1 for overview).
Equation (1) shows how different stimuli $S$ through weights $W$ affects how much the mood changes each time step ($\mathfrak{M}_{\delta}$), the $i$ is the index over the different sensors and their associated weights:
@@
Figure 1: An evolvable model of mood mechanism. Where different stimuli ($S$) and how they affect MoodMood (fluctuating affective state) and how mood affects behaviour, with weights ($W$) affecting the direction and strength of these effects. Therefore the weights define agents’ mood mechanism. Note: the dotted lines between $S_{2}$ and $S_{n}$ implies that there are more stimuli that can be in-between $S_{2}$ and $S_{n}$.
@@
Reproduction threshold The amount of energy an agent needs to have to reproduce. 10 Chosen for population stability, i.e. when its too low population regulation is very difficult as agents reproduce too much.
Agent metabolism Amount of energy agents consumes per time step not taking into account punishment. 0.1 Chosen so that amount of resource in environment is scarce enough to introduce a population regulation problem.
@@
Average growth rate Average replenishment of resource. 10 Chosen so that amount of resource in environment is scarce enough to introduce a population regulation problem.
@@ 3.9 Population level metrics
  • Hunger: the population average of how much of each agents bite size (how much you desire to eat) is unfulfilled $Hunger:=max(0,\mu_{eat}-resource)$.
  • Injuries from sanction: average energy loss across the population was just from being damaged by sanctions.
@@ 3.10 Notes on evolution and model of mood
Our model of evolution is not a complete model of evolution but merely attempts to show how optimisation propagates the type strategies that climb the adaptive gradient.
The model lends itself both cultural and genetic interpretations, since more successful strategies (attaining more resource) become more common in the population. In the genetic interpretation, genes determining behaviour are passed down with mutation and are more likely to be passed down if they are successful at attaining a resource. In the cultural interpretation it can be seen as strategies that are passed down culturally with a transmission error, with more successful strategies being imitated more to become the most common. The cultural interpretation allows for us to investigate the effect of normative punishment on norm emergence and the genetic interpretation which allows us to examine the effect of normative punishment on the evolution of mood. Further, this is a space-less simulation with no topology. This was done in order to focus on the effects normative punishment on the evolution of mood and not have it complicated by network topologies. Further we emphasise that the weights of the mood mechanism are being evolved, not the behaviours: $\mu_{eat}$ and $\mu_{sanction}$. They are produced by the evolved mood mechanism.
@@ 4 Results
In this section we will reiterate our definition of a axiomatic normative regularity and operationalise it in our simulation. We will compare two conditions:
@@
We asses whether punishment does indeed produce multiple equilibria and therefore differentiate it from other social processes.
@@
    The behaviour the population stabilises at is arbitrary across runs to a certain extent i.e. multiple equilibria across runs are reached : This criterion ensures the resultant behaviour is not fully due to environmental or social scaffolding, meaning when the behaviour is the only rational/viable action given the environmental constraints [48]. This would be indicated by repeated simulations stabilising at different average values i.e. there is a high variance of the means of each population.
Now that we have articulated our definition we will use it to interpret our simulation results, to confirm if explicit punishment processes and implicit ones lead to the dynamics predicted by the axiomatic normativity. In the explicit punishment condition, agents can punish each other according to their mood modulated behaviour $\mu_{sanction}$ which determines what $\mu_{eat}$ from another agent triggers a punishment. In the competition condition: Agents only have $\mu_{eat}$ and socially interact indirectly through the shared resource.
Overall, we see in Table 2 that punishment condition satisfies both conditions 1 and 2 of our definition. This is since we see a reduction in average variance of behaviour, signalling convergence, but still have variance between means, signalling arbitrariness. Therefore punishment results in behavioural regularities qualifying as axiomatically normative. Whereas implicit punishment i.e. competition satisfies only condition 1. convergence not 2. arbitrariness, therefore not qualifying as an axiomatically normative process. We can conclude our definition works for distinguishing between these two processes in simulation (see also Figure2).
Table 2: Variance of mean (extent of multi equilibria) and mean of variances (extent of behavioural convergence) for both the normative punishment and competition condition after 2000 simulation steps. The start of simulation values are also displayed to provide a comparison. We see that for both competition and punishment conditions conditions, that the mean of the variance of $eat_{\mu}$ decreases, thereby satisfying criterion 1 of our normative regularity definition. However, when we look at variance of the mean, we see that it almost zero in the competition condition (as individuals all converge to high $eat_{\mu}$. However in the punishment condition, we see much larger levels of variance across the means (an order of magnitude more), therefore maintaining a certain degree of arbitrariness. Therefore, only the punishment condition results in a behavioural regularities that meets both our criteria for being a normative regularity.
@@
Strong selection for $\mu_{eat}$ (how much resource the agent takes from environment), which leads to low variances in the means in both competition and punishment conditions Table 2. However, there is more variance of mean for $\mu_{eat}$ in the punishment condition. This suggests that this is not merely environmental scaffolding but a true normative regularity. Whereas the variance of mean for $\mu_{eat}$ in the non-minimally normative condition is almost 0 suggesting that even though there is a reduction in variance it is mostly down to environmental and social scaffolding and not truly normative. However, one could argue that both these behaviours are socially maintained (as per [48]); in the competition case agents compete with each other for resources and therefore implicitly regulate each other’s behaviour through leaving each other less resources in in the environment. So you could argue that emergent behaviours constitute "normative regularities" based on the definition of [48]. However our definition allows distinction between regularities based on whether the system tends to multiple equilibria or not. The key difference, between punishment and competition, is that the former allows for many equilibria but the other tends to one equilibrium: i.e. maximal $\mu_{eat}$, since that is the only direction of change that is permitted given the competition of the other agents, the system will tend to everyone being maximally selfish. So even though competition indirectly influences others behaviour through competition for resources in the environment, its only through explicit punishment, when the agents are directly effecting each others actions based on $\mu_{eat}$, that one gets many equilibria (as indicated by the higher variance between populations). This is due to the tension between punishment (explicit regulation of another’s cultural or biological fitness) and the drive to maximise utility, where many possible outcomes exist, as there are many ways one can balance these two pressures i.e. one can choose to take the punishment and then eat a lot to counteract it or one can choose to eat a little and not get punished, and then there are many such combinations in-between that are viable. Thus, through normative punishment there are now many possible equilibria that can be reached due this push and pull, which is more like true social norms in that they are determined by the group and somewhat arbitrary [4, 35] vs non normative social processes like competition that tend to one equilibrium.
@@
Figure 2: This figure depicts the average $\mu_{eat}$ for different simulation runs over time, with the black line being the average across runs and each coloured line the average of a population within a run. Comparing the competition and punishment condition, we see that competition tends to one equilibrium (the maximum value) whereas the punishment condition results in a wide range of equilibria that cover the whole spectrum of $\mu_{eat}$. (a) Competition: In brief, in competition there is a positive correlation between avg-$W_{hunger}$ (bad mood) and avg-$W_{bite}$ indicating a "bad mood → consume more" response (b) Punishment (axiomatic normativity): In brief, there is a positive correlation between avg-$W_{energy-gain}$ (good mood) and avg-$W_{bite}$ indicating a "bad mood → consume less" mechanism. Additionally, there is a positive correlation avg-$W_{others-being-punished}$ vs. avg-$W_{bite}$ indicating that others will lower their consumption when they see others being punished.
@@ 4.2 Different (social) mood mechanism emerges under punishment vs competition
Under non-normative social process(competition) vs. normative process (punishment) we see two different versions of mood mechanism emerge which rely on different stimuli to affect behaviour in different ways (Figure 3). The correlation matrix illustrates how the stimuli weights and behaviour weights are wired and therefore how a given stimulus will effect mood and subsequently how the mood will affect the behaviour.
@@
We see that in both conditions, information about energy (being taken in i.e. as signalled by hunger or energy-gain) is used. Thereby confirming, something not entirely surprising, that using the energy gained in the last state is useful for surviving.
@@
We see in the competition case (non normative) that we have a strong positive correlation with the avg-$W_{hunger}$ in relation to the avg-$W_{bite}$ (Figure 5 c) and weak positive weights with respect to avg-$W_{energy-gain}$ vs avg-$W_{bite}$ and avg-$W_{hunger}$ vs. avg-$W_{energy-gain}$ (Figure 5 a,b). This means that agents, when they lose energy, tend to consume more. This losing energy state could be argued to be negative mood since losing energy tracks with being in an unfavourable state  [^1] We have not pre-defined what a “good” mood is. Since both the input weights and output weights can evolve to values from -1.0 to 1.0 the system has one functional symmetry, i.e. we could invert all input and all output weights and would get the same functionality. As a result, there is no ”a priori” definition of what a ”good” mood is. For later interventions, we consider the good mood direction to be the one that correlates with positive energy gain as an input. The reason we do this is that the individuals that managed to survive in the simulation all had mood mechanisms that made them behave in a way to gain energy rather than lose it, otherwise they would not have survived. This also fits well with mood definitions that generally define it as “how well an organism is doing” [15] but again, we have let this arise through evolution and not hand coded it ourselves.. The mechanism we get from the population with competition is in line with what we would expect from a homeostatic model of affect: in response to negative mood, exert a large response to shift the system back into a favourable state (to not hungry in our case). This is similar to states like hunger or stress in the literature that use negative feedback like stress hormone cortisol [38, 29].
@@
We see that this is not the case with the punishment condition (normative case), where we have no correlation for avg-$W_{bite}$ vs. avg-$W_{hunger}$ and for avg-$W_{energy-gain}$ vs. avg-$W_{hunger}$ (Figure 4, b, c). Instead we have a strong positive correlation for avg-$W_{energy-gain}$ vs avg-$W_{bite}$ (Figure 4, a) meaning agents eat more when they are gaining energy (arguably like a good mood), and they eat less when losing energy (arguably a bad mood). This behaviour seems irrational as agents will not increase their consumption when they are losing energy and are, therefore, in a more precarious state. However, if we look at this behaviour in the context of punishment, losing energy is also correlated with being punished. Therefore in the context of punishment, it is rational for the system to lower its eating amount to avoid punishment and to increase it in the absence of punishment. This is further strengthened by the fact that there is a strong negative correlation between the avg-$W_{others-being-punished}$ vs. avg-$W_{bite}$, meaning agents: When they see others being punished, may indicate to them that they may be punished as well, so in response they reduce the amount they are eating to avoid punishment. Another observation worth mentioning is the clustering of values away from 0 in a and d, and to a lesser extent e and f. It seems unlikely for a population to not arrive at a relation between these weights, suggesting that there is an adaptive advantage to having this relation vs. not having it.
@@
This behaviour, eating more when an agent is a good mood and less when the agent is in a bad mood is the emergent encoding of a social preference, which takes into account the social consequences of taking too much resource. This social preference is encoded in how mood mechanism works itself, with punishment rewiring what it means to be in a "good" or "bad" mood. We emphasise that this emerges through evolution and punishment alone, without us hand coding the effects of stimuli on mood and mood on behaviour. Moreover, eating more in a favourable state (good mood) is similar to positive feedback based hormonal mechanisms e.g. social hormones like oxytocin which are involved in social and non social functions [41, 29]. Furthermore, the agents withdrawing from taking from the shared resource when in a "bad" mood, in order to avoid punishment, can be interpreted as a depressive mood. According to the social risk hypothesis, depression represents an adaptive response to the perceived threat of exclusion from important social relationships that, over the course of evolution, have been critical to maintaining an individual’s fitness prospects [3, 2].
@@
Further, there is a negative correlation between avg-$W_{energy-gain}$ vs. avg-$W_{sanction}$. This means that agents seem to punish less when they are losing energy and they are more tolerant to deviance when they are in a "bad mood" (they are losing energy either because there is no food/they are punished) and punish more when they are losing energy (bad mood). This seems irrational, but again if we look at this behaviour in the context of punishment, it allows for agents to minimise the punishment expenditure on energy when they are in a bad mood (they are losing energy either because there is no food/they are being punished) and therefore more efficient in a social environment [22].
@@
Figure 3: Here we see the effect of evolution on the correlations between different stimuli weights, behavioural weights which encode mood mechanism i.e. how stimuli affect mood and how mood affects behaviour respectively )as well as other evolvable agent traits) and how they are correlated with one another across populations after evolution. (a) competition (b) punishment (axiomatic normativity)
Figure 4: This figure depicts various averages of different traits plotted against each other for the condition with punishment (normative) to see the relation between traits over different runs. Each blue circle is a different simulation run and the blue line indicates the general trend. Here we see that avg-$W_{energy-gain}$ vs avg-$W_{bite}$ is the strongest positive correlation, indicating on average that agents will eat more if they are gaining energy and east less if they are losing energy. Further we see that there’s a negative correlation between avg-$W_{sanction}$ vs avg-$W_{energy-gain}$, indicating that agents will punish more (since sanction threshold is lower) when they are gaining energy. We also see negative correlation for avg-$W_{others-being-punished}$ vs avg-$W_{bite}$, indicating agents will eat less if they see others being punished. (a) avg-$W_{energy-gain}$ vs avg-$W_{bite}$ (b) avg-$W_{hunger}$ vs avg-$W_{energy-gain}$ (c) avg-$W_{hunger}$ vs avg-$W_{bite}$ (d) avg-$W_{others-being-punished}$ vs avg-$W_{bite}$ (e) avg-$W_{sanction}$ vs avg-$W_{energy-gain}$ (f) avg-$W_{sanction}$ vs avg-$W_{bite}$
Figure 5: This figure depicts various averages of different traits plotted against each other for the condition with competition (non-normative) to see the relation between traits over different runs. Each blue circle is a different simulation run and the blue line indicates the general trend. Here we see, in contrast to the punishment condition, a strong correlation between avg-$W_{hunger}$ vs avg-$W_{bite}$ and weak correlations for avg-$W_{energy-gain}$ vs avg-$W_{bite}$ and avg-$W_{hunger}$ vs avg-$W_{energy-gain}$. This means that on average agents will eat more if they are in a hungry state and losing energy (a) avg-$W_{energy-gain}$ vs avg-$W_{bite}$(b) avg-$W_{hunger}$ vs avg-$W_{energy-gain}$ (c) avg-$W_{hunger}$ vs avg-$W_{bite}$

### v4

@@ 3.8 Model of mood: an affective layer to modulate behaviour
We use mood mechanism to mean the tendency of an individual toward something, i.e. psychological inclination/disposition, and it can mean the environment’s power of disposing the individual to certain modes of interaction. Inspired by this, we define how mood is affected by stimuli (environment) and how mood, in turn, affects behaviour as a mood mechanism. We take inspiration for our model of mood from the empirically verified model of mood [14, 15]. More specifically, Eldar et al describe mood state as a scalar summary of "how well things are going": an average of recent reward error (positive or negative) over recent experiences in their case. We define it as average energy gain or loss in our case (it is our model’s equivalent of reward, see $S_{energy-gain}$ stimuli). We choose this model as it is a reification, capturing the essence of mood cognition instead of focusing on any particular emotion in any particular organism. Those core features being: stimuli (both internal and external) influencing a fluctuating mood state that in turn modulates behavioural and perceptual tendencies. We also add other stimuli as those chosen in [15]. We chose these stimuli based on the sensory information available to the agents that we believe to be important regarding their survival in the simulated environment (both physical and social). For example, hunger, being punished, seeing others violating a norm, and seeing others die. This list is not exhaustive, and in theory you could add more (hence the dotted lines between $Stimulus_{2}$ and $Stimulus_{n}$). We randomly initialise the importance of these stimuli and the direction of their effect on mood as a way of seeing when evolution (if at all) utilises them. In other words, we let our evolutionary process specify the functional relationship between its senses and its mood and in turn, its mood and its behaviour. This functional specification is what we call a mood mechanism. This allows us to see different kinds of mood mechanisms emerging without presupposing a mechanism. This allows us to explore how different types of social processes e.g. normative process like punishment pattern mood, produce different mood mechanisms.
@@
Now that we have described the overview of how the mood model operates, we will describe it in detail. What follows is a formal description of how mood affects behaviour in our model, with the weights (mood mechanism) affecting the direction and strength of this effect. In more detail, we describe how mood is updated based on stimuli and their evolvable weights, how mood updates with time and how mood consequently affects behaviour also through evolvable weights (see Figure 1 for overview).
Equation (1) shows how different stimuli $S$ through weights $W$ affect how much the mood changes each time step ($\mathfrak{M}_{\delta}$), the $i$ is the index over the different sensors and their associated weights:
@@
Figure 1: An evolvable model of mood mechanism. Where different stimuli ($S$) and how they affect MoodMood (fluctuating affective state), and how mood affects behaviour, with weights ($W$) affecting the direction and strength of these effects. Therefore, the weights define agents’ mood mechanism. Note: the dotted lines between $S_{2}$ and $S_{n}$ imply that there are more stimuli that can be in-between $S_{2}$ and $S_{n}$.
@@
Reproduction threshold The amount of energy an agent needs to have to reproduce. 10 Chosen for population stability, i.e. when it is too low, population regulation is very difficult as agents reproduce too much.
Agent metabolism Amount of energy agents consume per time step, not taking into account punishment. 0.1 Chosen so that the amount of resource in the environment is scarce enough to introduce a population regulation problem.
@@
Average growth rate Average replenishment of resource. 10 Chosen so that the amount of resource in the environment is scarce enough to introduce a population regulation problem.
@@ 3.9 Population level metrics
  • Hunger: the population average of how much of each agent’s bite size (how much you desire to eat) is unfulfilled $Hunger:=max(0,\mu_{eat}-resource)$.
  • Injuries from sanctions: average energy loss across the population was just from being damaged by sanctions.
@@ 3.10 Notes on evolution and model of mood
Our model of evolution is not a complete model of evolution, but merely attempts to show how optimisation propagates the type strategies that climb the adaptive gradient.
The model lends itself to both cultural and genetic interpretations, since more successful strategies (attaining more resources) become more common in the population. In the genetic interpretation, genes determining behaviour are passed down with mutation and are more likely to be passed down if they are successful at attaining a resource. In the cultural interpretation, it can be seen as strategies that are passed down culturally with a transmission error, with more successful strategies being imitated more to become the most common. The cultural interpretation allows for us to investigate the effect of normative punishment on norm emergence, and the genetic interpretation, which allows us to examine the effect of normative punishment on the evolution of mood. Further, this is a spaceless simulation with no topology. This was done in order to focus on the effects of normative punishment on the evolution of mood and not have it complicated by network topologies. Further, we emphasise that the weights of the mood mechanism are being evolved, not the behaviours: $\mu_{eat}$ and $\mu_{sanction}$. They are produced by the evolved mood mechanism.
@@ 4 Results
In this section, we will reiterate our definition of an axiomatic normative regularity and operationalise it in our simulation. We will compare two conditions:
@@
We assess whether punishment does indeed produce multiple equilibria and therefore differentiate it from other social processes.
@@
    The behaviour the population stabilises at is arbitrary across runs to a certain extent, i.e. multiple equilibria across runs are reached: This criterion ensures the resultant behaviour is not fully due to environmental or social scaffolding, meaning when the behaviour is the only rational/viable action given the environmental constraints [48]. This would be indicated by repeated simulations stabilising at different average values i.e. there is a high variance of the means of each population.
Now that we have articulated our definition, we will use it to interpret our simulation results, to confirm if explicit punishment processes and implicit ones lead to the dynamics predicted by the axiomatic normativity. In the explicit punishment condition: Agents can punish each other according to their mood modulated behaviour $\mu_{sanction}$, which determines what $\mu_{eat}$ from another agent triggers a punishment. In the competition condition: Agents only have $\mu_{eat}$ and socially interact indirectly through the shared resource.
Overall, we see in Table 2 that the punishment condition satisfies both conditions 1 and 2 of our definition. This is because we see a reduction in average variance of behaviour, signalling convergence, but still have variance between means, signalling arbitrariness. Therefore, punishment results in behavioural regularities qualifying as axiomatically normative. Whereas implicit punishment i.e. competition, satisfies only condition 1. convergence not 2. arbitrariness, therefore not qualifying as an axiomatically normative process. We can conclude our definition works for distinguishing between these two processes in simulation (see also Figure2).
Table 2: Variance of mean (extent of multi equilibria) and mean of variances (extent of behavioural convergence) for both the normative punishment and competition condition after 2000 simulation steps. The start of simulation values are also displayed to provide a comparison. We see that for both competition and punishment conditions, the mean of the variance of $eat_{\mu}$ decreases, thereby satisfying criterion 1 of our normative regularity definition. However, when we look at the variance of the mean, we see that it is almost zero in the competition condition (as individuals all converge to high $eat_{\mu}$. However, in the punishment condition, we see much larger levels of variance across the means (an order of magnitude more), therefore maintaining a certain degree of arbitrariness. Therefore, only the punishment condition results in a behavioural regularity that meets both our criteria for being a normative regularity.
@@
Strong selection for $\mu_{eat}$ (how much resource the agent takes from the environment), which leads to low variances in the means in both competition and punishment conditions, Table 2. However, there is more variance of the mean for $\mu_{eat}$ in the punishment condition. This suggests that this is not merely environmental scaffolding but a true normative regularity. Whereas the variance of mean for $\mu_{eat}$ in the non-minimally normative condition is almost 0 suggesting that even though there is a reduction in variance, it is mostly down to environmental and social scaffolding and not truly normative. However, one could argue that both these behaviours are socially maintained (as per [48]); in the competition case, agents compete with each other for resources and therefore implicitly regulate each other’s behaviour through leaving each other less resources in the environment. So you could argue that emergent behaviours constitute "normative regularities" based on the definition of [48]. However, our definition allows distinction between regularities based on whether the system tends to multiple equilibria or not. The key difference between punishment and competition is that the former allows for many equilibria, but the latter tends to one equilibrium: i.e. maximal $\mu_{eat}$, since that is the only direction of change that is permitted given the competition of the other agents, the system will tend to everyone being maximally selfish. So even though competition indirectly influences others’ behaviour through competition for resources in the environment, it’s only through explicit punishment, when the agents are directly affecting each other’s actions based on $\mu_{eat}$, that one gets many equilibria (as indicated by the higher variance between populations). This is due to the tension between punishment (explicit regulation of another’s cultural or biological fitness) and the drive to maximise utility, where many possible outcomes exist, as there are many ways one can balance these two pressures i.e. one can choose to take the punishment and then eat a lot to counteract it or one can choose to eat a little and not get punished, and then there are many such combinations in-between that are viable. Thus, through normative punishment, there are now many possible equilibria that can be reached due to this push and pull, which is more like true social norms in that they are determined by the group and somewhat arbitrary [4, 35] vs non-normative social processes like competition that tend to one equilibrium.
@@
Figure 2: This figure depicts the average $\mu_{eat}$ for different simulation runs over time, with the black line being the average across runs and each coloured line the average of a population within a run. Comparing the competition and punishment conditions, we see that competition tends to one equilibrium (the maximum value), whereas the punishment condition results in a wide range of equilibria that cover the whole spectrum of $\mu_{eat}$. (a) Competition: In brief, in competition there is a positive correlation between avg-$W_{hunger}$ (bad mood) and avg-$W_{bite}$ indicating a "bad mood → consume more" response (b) Punishment (axiomatic normativity): In brief, there is a positive correlation between avg-$W_{energy-gain}$ (good mood) and avg-$W_{bite}$ indicating a "bad mood → consume less" mechanism. Additionally, there is a positive correlation avg-$W_{others-being-punished}$ vs. avg-$W_{bite}$, indicating that others will lower their consumption when they see others being punished.
@@ 4.2 Different (social) mood mechanism emerges under punishment vs competition
Under non-normative social process(competition) vs. normative process (punishment), we see two different versions of mood mechanism emerge, which rely on different stimuli to affect behaviour in different ways (Figure 3). The correlation matrix illustrates how the stimuli weights and behaviour weights are wired and therefore how a given stimulus will affect mood and subsequently how the mood will affect the behaviour.
@@
We see that in both conditions, information about energy (being taken in, i.e. as signalled by hunger or energy-gain) is used. Thereby confirming, something not entirely surprising, that using the energy gained in the last state is useful for surviving.
@@
We see in the competition case (non normative) that we have a strong positive correlation with the avg-$W_{hunger}$ in relation to the avg-$W_{bite}$ (Figure 5 c) and weak positive weights with respect to avg-$W_{energy-gain}$ vs avg-$W_{bite}$ and avg-$W_{hunger}$ vs. avg-$W_{energy-gain}$ (Figure 5 a,b). This means that agents, when they lose energy, tend to consume more. This losing energy state could be argued to be negative mood since losing energy tracks with being in an unfavourable state.  [^1] We have not pre-defined what a “good” mood is. Since both the input weights and output weights can evolve to values from -1.0 to 1.0, the system has one functional symmetry, i.e. we could invert all input and all output weights and would get the same functionality. As a result, there is no ”a priori” definition of what a ”good” mood is. For later interventions, we consider the good mood direction to be the one that correlates with positive energy gain as an input. The reason we do this is that the individuals who managed to survive in the simulation all had mood mechanisms that made them behave in a way to gain energy rather than lose it; otherwise they would not have survived. This also fits well with mood definitions that generally define it as “how well an organism is doing” [15], but again, we have let this arise through evolution and not hand-code it ourselves. The mechanism we get from the population with competition is in line with what we would expect from a homeostatic model of affect: in response to negative mood, exert a large response to shift the system back into a favourable state (to not be hungry in our case). This is similar to states like hunger or stress in the literature that use negative feedback like stress hormone cortisol [38, 29].
@@
We see that this is not the case with the punishment condition (normative case), where we have no correlation for avg-$W_{bite}$ vs. avg-$W_{hunger}$ and for avg-$W_{energy-gain}$ vs. avg-$W_{hunger}$ (Figure 4, b, c). Instead, we have a strong positive correlation for avg-$W_{energy-gain}$ vs avg-$W_{bite}$ (Figure 4, a), meaning agents eat more when they are gaining energy (arguably like a good mood), and they eat less when losing energy (arguably a bad mood). This behaviour seems irrational as agents will not increase their consumption when they are losing energy and are, therefore, in a more precarious state. However, if we look at this behaviour in the context of punishment, losing energy is also correlated with being punished. Therefore, in the context of punishment, it is rational for the system to lower its eating amount to avoid punishment and to increase it in the absence of punishment. This is further strengthened by the fact that there is a strong negative correlation between the avg-$W_{others-being-punished}$ vs. avg-$W_{bite}$, meaning agents, when they see others being punished, may indicate to them that they may be punished as well, so in response, they reduce the amount they are eating to avoid punishment. Another observation worth mentioning is the clustering of values away from 0 in a and d, and to a lesser extent e and f. It seems unlikely for a population to not arrive at a relation between these weights, suggesting that there is an adaptive advantage to having this relation vs. not having it.
@@
This behaviour, eating more when an agent is in a good mood and less when the agent is in a bad mood, is the emergent encoding of a social preference, which takes into account the social consequences of taking too much resource. This social preference is encoded in how the mood mechanism works itself, with punishment rewiring what it means to be in a "good" or "bad" mood. We emphasise that this emerges through evolution and punishment alone, without us hand-coding the effects of stimuli on mood and mood on behaviour. Moreover, eating more in a favourable state (good mood) is similar to positive feedback-based hormonal mechanisms, e.g. social hormones like oxytocin, which are involved in social and non-social functions [41, 29]. Furthermore, the agents withdrawing from taking from the shared resource when in a "bad" mood, in order to avoid punishment, can be interpreted as a depressive mood. According to the social risk hypothesis, depression represents an adaptive response to the perceived threat of exclusion from important social relationships that, over the course of evolution, have been critical to maintaining an individual’s fitness prospects [3, 2].
@@
Further, there is a negative correlation between avg-$W_{energy-gain}$ vs. avg-$W_{sanction}$. This means that agents seem to punish less when they are losing energy, and they are more tolerant to deviance when they are in a "bad mood" (they are losing energy either because there is no food/they are punished) and punish more when they are losing energy (bad mood). This seems irrational, but again, if we look at this behaviour in the context of punishment, it allows for agents to minimise the punishment expenditure on energy when they are in a bad mood (they are losing energy either because there is no food/they are being punished) and therefore more efficient in a social environment [22].
@@
Figure 3: Here we see the effect of evolution on the correlations between different stimuli weights, behavioural weights which encode mood mechanism, i.e. how stimuli affect mood and how mood affects behaviour respectively (as well as other evolvable agent traits) and how they are correlated with one another across populations after evolution. (a) competition (b) punishment (axiomatic normativity)
Figure 4: This figure depicts various averages of different traits plotted against each other for the condition with punishment (normative) to see the relation between traits over different runs. Each blue circle is a different simulation run, and the blue line indicates the general trend. Here we see that avg-$W_{energy-gain}$ vs avg-$W_{bite}$ is the strongest positive correlation, indicating on average that agents will eat more if they are gaining energy and eat less if they are losing energy. Further, we see that there’s a negative correlation between avg-$W_{sanction}$ vs avg-$W_{energy-gain}$, indicating that agents will punish more (since the sanction threshold is lower) when they are gaining energy. We also see negative correlation for avg-$W_{others-being-punished}$ vs avg-$W_{bite}$, indicating agents will eat less if they see others being punished. (a) avg-$W_{energy-gain}$ vs avg-$W_{bite}$ (b) avg-$W_{hunger}$ vs avg-$W_{energy-gain}$ (c) avg-$W_{hunger}$ vs avg-$W_{bite}$ (d) avg-$W_{others-being-punished}$ vs avg-$W_{bite}$ (e) avg-$W_{sanction}$ vs avg-$W_{energy-gain}$ (f) avg-$W_{sanction}$ vs avg-$W_{bite}$
Figure 5: This figure depicts various averages of different traits plotted against each other for the condition with competition (non-normative) to see the relation between traits over different runs. Each blue circle is a different simulation run, and the blue line indicates the general trend. Here we see, in contrast to the punishment condition, a strong correlation between avg-$W_{hunger}$ vs avg-$W_{bite}$ and weak correlations for avg-$W_{energy-gain}$ vs avg-$W_{bite}$ and avg-$W_{hunger}$ vs avg-$W_{energy-gain}$. This means that on average agents will eat more if they are in a hungry state and losing energy (a) avg-$W_{energy-gain}$ vs avg-$W_{bite}$(b) avg-$W_{hunger}$ vs avg-$W_{energy-gain}$ (c) avg-$W_{hunger}$ vs avg-$W_{bite}$